# Spark: A System for Scientifically Creative Idea Generation

### v1

@@ Abstract
Recently, large language models (LLMs) have shown promising abilities to generate novel research ideas in science, a direction which coincides with many foundational principles in computational creativity (CC). In light of these developments, we present an idea generation system named Spark that couples retrieval-augmented idea generation using LLMs with a reviewer model named Judge trained on 600K scientific reviews from OpenReview. Our work is both a system demonstration and intended to inspire other CC researchers to explore grounding the generation and evaluation of scientific ideas within foundational CC principles. To this end, we release the annotated dataset used to train Judge, inviting other researchers to explore the use of LLMs for idea generation and creative evaluations.
@@ Introduction
Addressing grounding explicitly, SciPIP integrated semantic retrieval and citation awareness into idea generation, producing hypotheses informed by current literature but primarily evaluated within NLP domains (?). Extending this further, SciMON employed iterative refinement to systematically maximize novelty, though practical feasibility and experimental utility remained secondary concerns (?).
@@ Methods
The Spark system integrates advanced AI components designed to support computational creativity in scientific idea generation, leveraging specialized agentic systems for literature retrieval, idea synthesis, refinement, and evaluation. Figure 1 depicts the integrated process of the Spark and Xplor systems.
@@
Xplor, inspired by PaperQA (?), is designed to address the specific literature retrieval and analysis needs of the Spark system. It employs sophisticated embedding techniques alongside maximal marginal relevance (MMR) for identifying, collecting, and analyzing research papers relevant to a given research question or topic (?).
@@
The foundation of Xplor is OpenAI’s text-embedding model, text-embedding-3-large (?), which transforms research papers into rich 1536-dimensional semantic vectors. These embeddings facilitate accurate semantic comparisons among documents. The embeddings are indexed using the FAISS library to allow efficient nearest-neighbor searches based on cosine similarity between query embeddings and document embeddings (?).
Maximal Marginal Relevance:
After initial retrieval, Xplor applies an MMR-based re-ranking process to optimize the selection of relevant yet diverse papers (?). This approach iteratively selects documents with high relevance while simultaneously reducing redundancy, ensuring an informative and varied literature corpus.
Recursive Query Generation:
To further refine search results, Xplor uses a recursive query generation module. In autonomous mode, Xplor initially analyzes retrieved content using an LLM employing chain-of-thought prompting to identify key themes and subtopics. This process generates increasingly targeted follow-up queries, continuing until predefined relevance metrics signal sufficient depth of literature. In interactive mode, this refinement is driven by direct user or AI feedback, enabling precise, iterative query optimization.
@@
Mitigating Hallucinations in AI-Driven Scientific Creativity through Citation Network Analysis and Knowledge Graph Validation This research proposes a method for mitigating hallucinations in Large Language Models (LLMs) used for creative scientific idea generation by leveraging citation-network analysis and knowledge graphs. The approach involves constructing a knowledge graph from scientific literature—using datasets such as the Semantic Scholar Open Research Corpus (S2ORC)—to represent concepts and relationships within a specific scientific domain. When an LLM generates a scientific idea, its cited references are analyzed within the context of this knowledge graph. Unusual combinations of cited references, identified through citation-network analysis, are flagged for further scrutiny. The content of the cited publications, processed using models like SciBERT, is then compared against the LLM’s generated statements. Discrepancies between the LLM’s output and the information contained in the cited references, or inconsistencies with the established knowledge graph, are flagged as potential hallucinations. However, if the unusual combination leads to a supported and insightful idea, it is considered a potentially creative contribution. This method aims to improve the reliability of AI-driven scientific discovery while exploring the possibility of leveraging controlled “hallucinations” as a source of creative insights. The research contributes to the debate surrounding AI creativity by providing a framework for evaluating the validity and potential of LLM-generated scientific ideas.
@@
Idea Generation with Spark Idea Generator
The Spark Idea Generator constitutes the initial creative stage of the Spark system, formulating seed research proposals grounded in rigorous prior work analysis. It synthesizes concepts and identifies research challenges to generate structured, contextually relevant ideas.
Concept Extraction and Problem Synthesis: Initially, the process involves analyzing a corpus of scientific papers to extract key terminological clusters and high-level concepts. Concurrently, Xplor recursively identifies emergent research gaps and themes through related papers. These outputs construct a problem-driven research prompt (Figure 2, A-B).
Structured Idea Formulation: Using the constructed prompt, an LLM agent generates a preliminary research proposal comprising a structured reasoning plan, novel concepts, and a concise title and abstract, guided to incorporate relevant domain knowledge and maintain logical coherence (Figure 2, C-D). This output (as illustrated in Figure 4) forms the basis for high-quality, grounded research ideas suitable for advanced development.
@@
Review Generation: In this phase, each idea—represented by its title and abstract—is processed by our supervised fine-tuned (SFT) Judge model trained on OpenReview peer review corpora. This model generates multiple critiques per idea, highlighting strengths, limitations, and areas for improvement, simulating diverse expert perspectives.
Decision Synthesis: Subsequently, a separate decision-making model integrates these critiques, rendering a final binary classification—ACCEPT or REJECT. The model aggregates the earlier critiques to produce a reasoned decision and utility score (Figure 3, A-D).
This pipeline has successfully produced and filtered over 10,000 AI-focused research ideas, demonstrating its scalability and effectiveness for large-scale computational creativity tasks. In future work, it is worth exploring techniques to enhance the creativity and diversity of Spark’s generated ideas (??).
@@ Discussion
Several design choices merit further exploration. First, the current filter operates only once per idea; future studies could examine closed-loop variants that feed reviewer feedback back into the generator. Second, while Xplor’s retrieval is similarity-based, structured knowledge sources (e.g. ontologies) might support deeper analogical reasoning. Third, the present peer-review simulation produces textual critiques but no explicit explanatory traces; richer explanation outputs could help users understand the basis of automatic judgments. Finally, Spark has been evaluated only in a controlled setting confined to idea generation within the field of AI, but to understand its abilities across other disciplines, studies in additional scientific fields are required.
To facilitate community experimentation, we are releasing the annotated OpenReview dataset used to train the Judge model. We hope this proof-of-concept encourages comparative studies of retrieval methods, generation prompts, reviewer-agent designs, and creativity evaluation strategies using our system as a guide.
@@ Conclusion
This paper presented Spark as a system demonstration that unifies a literature retrieval engine, an LLM-driven idea-generation module, and a peer-review-inspired filter. The implementation is offered as a concrete step towards grounding scientific idea generation in foundational computational creativity principles rather than as a definitive solution to open challenges in AI-driven discovery. By illustrating how creative generation and evaluation can be combined, we aim to inspire new directions in the computational creativity community around the generation, evaluation, and feasibility of scientific idea generation using LLM systems. We anticipate that the framework, data resources, and observations reported here will serve as a starting point for future work exploring the relationship between computational creativity and scientific idea generation.

### v2

@@ Abstract
Recently, large language models (LLMs) have shown promising abilities to generate novel research ideas in science, a direction which coincides with many foundational principles in computational creativity (CC). In light of these developments, we present an idea generation system named Spark that couples retrieval-augmented idea generation using LLMs with a reviewer model named Judge trained on 600K scientific reviews from OpenReview. Our work is both a system demonstration and intended to inspire other CC researchers to explore grounding the generation and evaluation of scientific ideas within foundational CC principles. To this end, we invite other researchers to explore the use of LLMs for idea generation and creative evaluations, and release the annotated dataset used to train Judge. [^1] https://huggingface.co/datasets/spiralworks/openreview_wildcard_2025_v2
@@ Introduction
Addressing grounding explicitly, SciPIP integrated semantic retrieval and citation awareness into idea generation, producing hypotheses informed by current literature but primarily evaluated within NLP domains (?). Extending this further, SciMON employed iterative refinement to systematically maximize the novelty of ideas, but their outputs often suffered from low practical feasibility and experimental utility (?).
@@ Methods
The Spark system integrates AI-driven components that support computational creativity in the generation of scientific ideas, utilizing specialized agent-based modules for the retrieval of literature and the synthesis, refinement and evaluation of ideas. Figure 1 illustrates the interaction between the Spark and Xplor systems.
@@
Inspired by PaperQA (?), Xplor handles literature retrieval and analysis tasks tailored specifically for the Spark system. It utilizes OpenAI embeddings, FAISS indexing, maximal marginal relevance (MMR), and a recursive, agent-based retrieval loop to systematically gather, analyze, and refine relevant research documents to later be used for idea generation.
@@
Xplor employs OpenAI’s text-embedding-3-large model (?) to convert retrieved documents into 1536-dimensional semantic vectors. These embeddings allow accurate semantic comparison of research papers. FAISS (?) efficiently indexes these vectors, enabling rapid retrieval through cosine similarity-based nearest-neighbor searches when provided with user-generated input queries.
Relevance Scoring and Maximal Marginal Relevance:
After an initial vector-based retrieval, each candidate text chunk is also evaluated using an LLM-based summary model. This model assigns relevance scores (1 to 10) to each chunk according to its relevance to the research query, complementing the earlier cosine similarity measures. Subsequently, Xplor applies maximal marginal relevance (MMR) to re-rank these scored chunks, a technique which prioritizes both relevance and diversity. This ensures that selected documents collectively provide comprehensive coverage with minimal redundancy.
Recursive Retrieval and Evidence Gathering Loop:
The retrieval process in Xplor is recursive and agent-driven. Initially, the system queries databases (e.g., arXiv, Google Scholar) using a programmable search engine, retrieving documents relevant to the initial question. First, documents are segmented into overlapping chunks and embedded. Then, an agentic loop iteratively refines the search by generating new, increasingly precise queries informed by intermediate summaries from an LLM response. This process continues until specific depth and coverage criteria, such as accumulating at least five evidence snippets from multiple sources, are met.
@@
Xplor has two distinct operating modes: interactive and autonomous. In interactive mode, Xplor is directly prompted by the user query, whereas in autonomous mode, Xplor takes queries from the Spark Idea Generator agent to aid in the generation and refinement of ideas, as explained in the following section.
Idea Generation with Spark Idea Generator
The Spark Idea Generator constitutes the initial creative stage of the Spark system, formulating seed research proposals based on the earlier Xplor search results. The idea generation system identifies open research problems related to the query and then synthesizes concepts to generate structured, contextually relevant, and well-motivated ideas.
@@
Entropy Reduction and Intrinsic Motivation in AI Music Creation This research explores the simulation of human music creation using AI, focusing on incorporating intrinsic motivation and addressing the potential impact on human artists. We propose a system based on a Transformer network trained with reinforcement learning to generate music. The core idea is to model musical creativity as a process of ’restructuring high psychological entropy material,’ where entropy represents the complexity and unpredictability of musical elements. The AI is trained with an intrinsic reward function that encourages it to reduce the entropy of its generated music, quantified using Kullback-Leibler Divergence between probability distributions of musical features before and after transformations. This reward is shaped to encourage exploration and discovery of new musical structures. The system’s performance will be evaluated by comparing its generated music to human-composed music using perceptual audio features (MFCCs, chromagrams) and subjective ratings through metrics like cosine similarity and t-SNE visualization. Furthermore, we will conduct a survey to assess how exposure to the AI-generated music affects human artists’ motivation and creative output, measuring changes in their creative output and self-reported motivation levels. This research aims to provide insights into the computational modeling of creativity and address the ethical and societal implications of AI in the arts. It seeks to understand how AI can be a tool that both generates and inspires music.
@@
Concept Extraction and Problem Synthesis: Initially, a user provided research question guides Xplor to recursively retrieve relevant papers and identify open research problems. The concepts and open problems from these papers are then used to construct a problem-driven research prompt (Figure 2, A-B).
Structured Idea Formulation: Using the constructed prompt, an LLM agent generates a preliminary research proposal comprising a structured reasoning plan, novel concepts, and a concise title and abstract, guided to incorporate relevant domain knowledge and maintain logical coherence (Figure 2, C-D). This output (as illustrated in Figure 4) forms the basis for high-quality research ideas already contextualized within existing literature and ready to be explored by users of the system. If further refinement is needed, the system adaptively uses Xplor to retrieve more relevant literature to address a specific area of improvement.
@@
Review Generation: In this phase, each idea—represented by its title and abstract—is processed by our supervised fine-tuned (SFT) Judge model trained on OpenReview peer review corpora. This model generates multiple critiques per idea, highlighting strengths, limitations, and areas for improvement, simulating the diverse perspectives that might be offered during the scientific peer review process.
Decision Synthesis: Subsequently, a separate LLM integrates these reviews to produce a final binary classification—ACCEPT or REJECT—of the idea, an explanation of the decision, and utility score for the idea (Figure 3, A-D).
This pipeline has successfully produced and filtered over 10,000 AI-focused research ideas, demonstrating its scalability for large-scale computational creativity tasks. In future work, it is worth exploring techniques to enhance the creativity and diversity of Spark’s generated ideas (??).
@@
Nevertheless, several design choices merit further exploration and refinement. First, the current filter operates only once per idea; future studies could examine closed-loop variants that feed reviewer feedback back into the generator. Second, while Xplor’s retrieval is similarity-based, structured knowledge sources (e.g. ontologies) might support deeper analogical reasoning. Third, the present peer-review simulation produces textual critiques but no explicit explanatory traces; richer explanation outputs could help users understand the basis of automatic judgments. Finally, Spark has been evaluated only in a controlled setting confined to idea generation within the field of AI, but to understand its abilities across other disciplines, studies in additional scientific fields are required.
In summary, we hope this proof-of-concept can serve as a guide for future work in this area, encouraging comparative studies of retrieval methods, generation prompts, reviewer-agent designs, and creativity evaluation strategies using our system as a foundation. To facilitate community experimentation, we have released the annotated dataset, consisting of papers and reviews, used to train the Judge model.
Looking Ahead: Computational Creativity and Scientific Idea Generation
The use of AI systems to generate scientific ideas raises important considerations that the CC community is well-poised to address. We state some open questions on this topic below, with the hope of inspiring other CC researchers to begin exploring these issues in future work.
Creative Criteria for Ideas Versus Papers
Creativity is typically measured by three dimensions: novelty, utility, and surprise (???). While the utility of a scientific paper is mostly judged by its future impact—whether that be citation count, accolades like best paper awards, and other retrospective metrics—idea generation and evaluation systems such as ours are tasked with finding high-quality directions that merit further exploration before their research outcomes are ever observed. How should CC researchers approach utility evaluations for scientific ideas? We believe a combination of theory-guided and data-driven approaches may be most fruitful here, with our dataset release helping facilitate the latter.
Transformative Idea Generation
Similarly, creative processes are typically divided into three types: combinatorial, exploratory, and transformational (?). Our idea generation system relies on concept extraction and problem synthesis to construct well-motivated and contextually relevant ideas, often through the combination of existing concepts, as outlined in Figure 2. Nevertheless, many historically groundbreaking discoveries have been transformational, actively restructuring elements of the conceptual spaces afforded by their scientific field. Can AI systems be used to discover transformational scientific ideas? Considering the historical impact transformational creativity has had on science and society at large, we believe this direction merits considerable future exploration.
@@ Conclusion
This paper presented Spark as a system demonstration that unifies a literature retrieval engine, an LLM-driven idea-generation module, and a peer-review-inspired filter. The implementation and discussion in this work are offered as a concrete step towards grounding scientific idea generation in foundational computational creativity principles. By illustrating how creative generation and evaluation can be combined, we aim to inspire new directions in the computational creativity community around the generation, evaluation, and feasibility of scientific idea generation using LLM systems. We anticipate that the framework, data resources, and observations reported here will serve as a starting point for future work exploring the relationship between computational creativity and scientific idea generation.